John is evaluating a stock that he believes will pay a dividend equal to $2.75 in exactly one year from today. Further, John expects that in addition to the dividend he will be able to sell the stock for the amount $45 per share also one year from today. Using this information, what is the stock's price today if the required rate of return on the stock is 11.00%?
Answer #1

Answer:

Cash flows after one year from today = Dividend + Sale value of share = 2.75 + 45 = $47.75

Required rate of return on the stock is = 11.00%

Stock's price today = 47.75 / (1 + 11%) = $43.02

Stock's price today = $43.02